### 13 Mar 2024 v1

# The $q$-ary Gilbert–Varshamov bound can be improved for all but finitely many positive integers $q$

Abstract
For any positive integer $q\geq 2$ and any real number $\delta\in(0,1)$, let $\alpha_{q}(n,\delta n)$ denote the maximum size of a subset of $\mathbb{Z}_{q}^{n}$ with minimum Hamming distance at least $\delta n$, where $\mathbb{Z}_{q}=\{0,1,\dotsc,q-1\}$ and $n\in\mathbb{N}$. The asymptotic rate function is defined by $R_{q}(\delta)=\limsup_{n\rightarrow\infty}\frac{1}{n}\log_{q}\alpha_{q}(n,\delta n).$ The famous $q$-ary asymptotic Gilbert–Varshamov bound, obtained in the 1950s, states that
$R_{q}(\delta)\geq 1-\delta\log_{q}(q-1)-\delta\log_{q}\frac{1}{\delta}-(1-\delta)\log_{q}\frac{1}{1-\delta}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}R_{\mathrm{GV}}(\delta,q)$
for all positive integers $q\geq 2$ and $0<\delta<1-q^{-1}$. In the case that $q$ is an even power of a prime with $q\geq 49$, the $q$-ary Gilbert–Varshamov bound was firstly improved by using algebraic geometry codes in the works of Tsfasman, Vladut, and Zink and of Ihara in the 1980s. The further investigation in algebraic geometry codes has shown that the $q$-ary Gilbert–Varshamov bound can also be improved in the case that $q$ is an odd power of a prime but not a prime with $q>125$. However, it remains a long-standing open problem whether the $q$-ary Gilbert–Varshamov bound would be tight for those infinitely many integers $q$ which is a prime, except for Fermat primes not less than 257, and which is a generic positive integer not being a prime power.
In this paper, we prove that the $q$-ary Gilbert–Varshamov bound can be improved for all but finitely many positive integers $q\geq 2$. In particular, it is explicitly shown that $R_{q}(1/2)>R_{\mathrm{GV}}(1/2,q)$ for all integers $q>\exp(29)$. Furthermore, based on the $q$-ary asymptotic Plotkin upper bound $R_{q}(\delta)\leq 1-\delta-\frac{\delta}{q-1}$ for all positive integers $q\geq 2$ and $0<\delta<1-q^{-1}$, we can define the growth of the rate function $R_{q}(\delta)$ for $\delta\in(0,1)$ fixed and $q$ growing large as follows
$\eta(\delta)=\liminf_{q\rightarrow\infty}\frac{1}{\log q}\log[1-\delta-R_{q}(\delta)]^{-1}.$
The $q$-ary Plotkin upper bound implies that the growth satisfies $\eta(\delta)\leq 1$, while the $q$-ary Gilbert–Varshamov lower bound only provides the trivial bound $\eta(\delta)\geq 0$. In this paper, we show that the growth has actually a nontrivial lower bound as $\eta(\delta)\geq 1/6$ for $\delta\in(0,1)$. These new lower bounds, which exceed the Gilbert–Varshamov bound, are achieved by using codes from geometry of numbers introduced by Lenstra in the 1980s.
1 Introduction
We consider as an alphabet any finite set of $q$ symbols such as $\mathbb{Z}_{q}=\{0,1,\dotsc,q-1\}$ for any positive integer $q\geq 2$. Let $\mathbb{Z}_{q}^{n}$ denote the set of all $n$-tuples $x=(x_{1},\dotsc,x_{n})$ with components in $\mathbb{Z}_{q}$ for any positive integer $n\geq 1$. We define a metric on the set $\mathbb{Z}_{q}^{n}$ by letting $d_{\mathrm{H}}(x,y)$ be the number of positions in which the components of the two $n$-tuples $x$ and $y$ in $\mathbb{Z}_{q}^{n}$ differ, which is called the Hamming distance between $x$ and $y$. A $q$-ary code with length $n$ is a nonempty subset ${\cal C}$ of $\mathbb{Z}_{q}^{n}$ and its elements are called codewords. Let $\mathbb{N}$ and $\mathbb{P}$ denote the sets of positive integers and of prime numbers, respectively. If the size of ${\cal C}$, denoted by $|{\cal C}|$, is $M\in\mathbb{N}$ with $M\geq 2$ and the minimum Hamming distance between any two distinct codewords in ${\cal C}$ is $d$, then ${\cal C}$ is called a $q$-ary $(n,M,d)$ code. The rate $R$ of the code ${\cal C}\subseteq\mathbb{Z}_{q}^{n}$ is defined by $R=\frac{1}{n}\log_{q}|{\cal C}|,$ which is a value in $[0,1]$. For an arbitrary real number $\delta\in(0,1)$, let $\alpha_{q}(n,\delta n)$ denote the maximum size of a $q$-ary code with length $n$ which has the minimum Hamming distance at least $\delta n$. We define the asymptotic rate function as follows
$R_{q}(\delta)=\limsup_{n\rightarrow\infty}\frac{1}{n}\log_{q}\alpha_{q}(n,\delta n).$
The determination of the rate function $R_{q}(\delta)$ is an asymptotic sphere packing problem in the space $\mathbb{Z}_{q}^{n}$ with Hamming distance as $n\to\infty$, which has been considered as a central problem in combinatorial coding theory (see van Lint [45] and Manin [22]). The asymptotic rate function $R_{q}(\delta)$ plays a fundamental role in Shannon’s information theory as well (see [36] and [37]). It is intimately related to the achievable probability of decoding error in reliable transmission of information, which approaches 00 exponentially as the number $n$ of channel uses goes to infinity. The better a lower bound for the rate function $R_{q}(\delta)$ exists, the larger the error-exponent can be achieved by a sequence of error-correcting codes.
It is known that for all integers $q\geq 2$, if $11-q^{-1}\leq\delta<$ 1, then $R_{q}(\delta)=0$. The exact value of $R_{q}(\delta)$, however, is not known for any integer $q\geq 2$ and real number $\delta\in(0,1-q^{-1})$. A famous lower bound for $R_{q}(\delta)$ is the following $q$-ary asymptotic Gilbert–Varshamov bound [10], [46]
$R_{q}(\delta)\geq 1-\delta\log_{q}(q-1)-\delta\log_{q}\frac{1}{\delta}-(1-\delta)\log_{q}\frac{1}{1-\delta}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}R_{\mathrm{GV}}(\delta,q)$
for all integers $q\geq 2$ and $0<\delta<1-q^{-1}$.
Although it was obtained in the 1950s by simple combinatorial arguments in [10] or with the probabilistic method in [46], the $q$-ary asymptotic Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ was thought for a long time that it would be equal to the rate function $R_{q}(\delta)$.
Let $\mathbb{F}_{q}$ be a finite field with $q$ elements, where $q$ is a prime power. A $q$-ary code ${\cal C}\subseteq\mathbb{F}_{q}^{n}$ with length $n$ is called linear if it is a subspace of the vector space $\mathbb{F}_{q}^{n}$ over $\mathbb{F}_{q}$. In 1981, Goppa [12] found a remarkable connection between algebraic curves over $\mathbb{F}_{q}$ and $q$-ary linear codes and showed, through the Riemann–Roch theorem, that a better code can be obtained from an algebraic curve over $\mathbb{F}_{q}$ with more rational points given its genus. In 1982, by using Goppa codes [12] from algebraic curves over finite fields, Tsfasman, Vladut, and Zink [44] employed the classic modular curves and the Shimura curves, and successfully improved the Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ for $q$ being a prime power of $p^{2}$ ($p\geq 7$) and of $p^{4}$ ($p\geq 3$) and $\delta$ in an open subinterval of $(0,1-q^{-1})$. This discovery was the first breakthrough after over twenty years of searching for a lower bound to beat the Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ for certain $q$ and $\delta$.
We are going to present briefly theory of Goppa codes from algebraic curves over finite fields. The following work in this paper on lower bounds of the rate function will be based on Lenstra codes from geometry of numbers, which can be viewed as being analogous to Goppa codes.
In 1982, Manin [22] firstly described such connection for algebraic geometers with some arising unsolved problems for drawing their attention. In 1982, Ihara [16] defined the following number
$A(q)=\limsup_{g\rightarrow\infty}\frac{N_{q}(g)}{g}$ (1)
where $N_{q}(g)$ is the maximum number of rational points that a non-singular absolutely irreducible algebraic curve over $\mathbb{F}_{q}$ of genus $g$ can have. Ihara’s number $A(q)$ is well-defined by the Hasse–Weil theorem, namely, the Riemann hypothesis for algebraic curves over finite fields [40]. Ihara [16] obtained that $A(q)\geq\sqrt{q}-1$ when $q=p^{2m}$, $p$ prime and $m\in\mathbb{N}$, and provided an upper bound for $A(q)$ for all $q$ being a prime power. It has been proved that Goppa codes imply (see Manin [22, p. 718])
$R_{q}(\delta)\geq 1-\delta-\frac{1}{A(q)}$ (2)
for all $q$ being a prime power and $10<\delta<$ 1. According to Manin [22, p. 719], the work in Tsfasman, Vladut, and Zink [44], independently of the earlier Ihara’s work [14], [15] in the 1970s on towers of curves over $\mathbb{F}_{q}$, showed that $A(q)\geq\sqrt{q}-1$ when $q=p^{2}$ or $q=p^{4}$, $p$ prime, and hence from (2) that the Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ was improved for $q=p^{2}$ ($p\geq 7$) or $q=p^{4}$ ($p\geq 3$) and $\delta$ in an interval of $(\delta_{1},\delta_{2})$ including $\frac{q-1}{2q-1}$ (see [44, p. 28]). It is noted that a better upper bound of $A(q)$ was also obtained, which is called as the Drinfeld–Vladut bound [47], asserting that $A(q)\leq\sqrt{q}-1$ for all $q$ being a prime power. It had improved a previous upper bound of Ihara [16]. The Drinfeld–Vladut bound is tight when $q=p^{2m}$, $p$ prime and $m\in\mathbb{N}$, from the aforementioned lower bound of $A(q)$ due to Ihara [16].
Goppa codes from algebraic curves over finite fields are also called as algebraic geometry codes (see the monographs [42] and [43]). It is well-known that algebraic curves over $\mathbb{F}_{q}$ can be equivalently replaced in the language of algebraic function fields over $\mathbb{F}_{q}$ [40]. In the case of $q$ being an even power of a prime, Garcia and Stichtenoth [9] found an alternative approach with the Artin–Schreier tower of algebraic function fields over $\mathbb{F}_{q}$ to attain the Drinfeld–Vladut bound. When $q$ is an odd power of a prime but not a prime with $q>125$, Bassa, Beelen, Garcia, and Stichtenoth [3] showed that the Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ can also be improved by using appropriate infinite towers of algebraic function fields over finite fields $\mathbb{F}_{q}$.
In 1983, Serre [33], [34] demonstrated that there exists a constant $c>0$ such that $A(q)\geq\frac{1}{c}\log q$ for any integer $q\geq 2$ being a prime power, by using class field theory for algebraic function fields over $\mathbb{F}_{q}$ (see also the recent monograph by Serre [35]). This bound remains the best known lower bound for $A(q)$ in general for $q$ being a prime. By using (2), we have the following Serre’s lower bound
$R_{q}(\delta)\geq 1-\delta-\frac{c}{\log q}$ (3)
for any integer $q\geq 2$ being a prime power and $0<\delta<1-q^{-1}$. It was shown in Niederreiter and Xing [27] that the constant $c$ in (3) can be chosen as $96\log 2$ for $q\geq 2$ being a prime power or $75\log 2$ for $q\geq 2^{24}$ being a prime power. With this explicit constant $c$, it can be checked that Serre’s lower bound (3) is not better than the Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ for any integer $q$ being a prime and $0<\delta<1-q^{-1}$.
For $q\geq 2$ being an even power of a prime, i.e., $q=p^{2m}$, $p$ prime and $m\in\mathbb{N}$, in 2001 Elkies [7] found approaches to extend a sequence of Goppa codes over $\mathbb{F}_{q}$, which exceeds the $q$-ary Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ for certain $\delta$, to a sequence of codes over an alphabet with $q+1$ symbols which is then better than the $(q+1)$-ary Gilbert–Varshamov bound for certain $\delta$. Note that when $p=2$, Fermat numbers $F_{n}=2^{2^{n}}+1$, $n\in\mathbb{N}$, are included in the form of $q+1=2^{2m}+1$ with $m=2^{n-1}$.
Despite the above considerable improvement by algebraic geometry codes from curves over finite fields, the $q$-ary Gilbert–Varshamov bound has still been, since the 1950s, the best known lower bound for any positive integer $q\geq 2$ which is a prime, except for Fermat primes not less than 257, and for a generic positive integer $q$ not being a prime power in which case the finite field $\mathbb{F}_{q}$ with $q$ elements does not exist. Specifically, there are infinitely many primes $q\geq 2$ and infinitely many integers $q\geq 2$ not being prime powers for which the $q$-ary Gilbert–Varshamov bound is the current largest known lower bound. It remains a long-standing open problem whether the $q$-ary Gilbert–Varshamov bound would be tight for those infinitely many integers $q$. The answer for this problem has been beyond the reach of algebraic geometry codes in the previous works.
In this paper, we are going to show that the $q$-ary Gilbert–Varshamov bound can be improved for all but finitely many positive integers $q\geq 2$. In particular, we will establish the following explicit result.
Theorem 1
For all integers $q>\exp(29)$, we have
$R_{q}(1/2)>R_{\mathrm{GV}}(1/2,q).$ (4)
Thus, the $q$-ary Gilbert–Varshamov bound is possibly tight only for some $q<\exp(29)$.
The rate function $R_{q}(\delta)$ has the following $q$-ary asymptotic Plotkin upper bound [28], [45, p. 68]
$R_{q}(\delta)\leq 1-\delta-\frac{\delta}{q-1}$ (5)
for all positive integers $q\geq 2$ and $0<\delta<1-q^{-1}$, which is the best known upper bound for large $q$. Therefore, we can define the growth of the rate function $R_{q}(\delta)$ for $\delta\in(0,1)$ fixed and $q$ growing large by
$\eta(\delta)=\liminf_{q\rightarrow\infty}\frac{1}{\log q}\log[1-\delta-R_{q}(\delta)]^{-1}.$ (6)
The $q$-ary Plotkin upper bound implies that the growth satisfies $\eta(\delta)\leq 1$. On the other hand, the $q$-ary asymptotic Gilbert–Varshamov bound can be represented by
$R_{\mathrm{GV}}(\delta,q)=1-\delta-\frac{h(\delta)}{\log q}+O\left(\frac{1}{q\log q}\right)$ (7)
as $q\rightarrow\infty$, where $h(\delta)=-\delta\log\delta-(1-\delta)\log(1-\delta),$ $10<\delta<$ 1, is the entropy function. The $q$-ary Gilbert–Varshamov lower bound only gives the trivial bound $\eta(\delta)\geq 0$. In this paper, we will show that the growth $\eta(\delta)$ has a nontrivial lower bound as follows.
Theorem 2
For all real numbers $\delta\in(0,1)$, we have
$\frac{1}{6}\leq\eta(\delta)\leq 1.$ (8)
It is noticed that the Gilbert–Varshamov bound in (7) indicates that the reciprocal of the difference between $1-\delta$ and $R_{q}(\delta)$ has asymptotically at least a logarithmic order in $q$. Theorem 8 demonstrates that the reciprocal of the difference between $1-\delta$ and $R_{q}(\delta)$ has asymptotically an algebraic order, rather than a logarithmic order, in $q$.
The new lower bounds in Theorem 1 and Theorem 8 are based on theory of Lenstra codes from geometry of numbers, which can be viewed as an analogue of theory of Goppa codes from algebraic curves over finite fields. In Section 2, we will present a detailed account of a variant of Lenstra codes. In Section 3, a theory of infinite towers of Hilbert class fields and its combination with Lenstra codes are developed. The theories presented in Section 2 and in Section 3 enable us to obtain an applicable lower bound for the rate function $R_{q}(\delta)$ in Section 4, which will be used in proofs of Theorem 1 and Theorem 8. With those preparations, we will establish proofs of Theorem 1 and Theorem 8 in Section 5 and Section 6, respectively.
2 Codes from geometry of numbers
Let $K$ be an algebraic number field with degree $m=[K:\mathbb{Q}]$ over the field $\mathbb{Q}$ of rational numbers. If $K$ has $s$ real embeddings and $t$ pairs of complex conjugate embeddings, then we have $s+2t=m$. We denote by $O_{K}$ the ring of integers of $K$ and by $\Delta_{K}$ the discriminant of $O_{K}$ over the ring $\mathbb{Z}$ of rational integers. In Minkowski’s theory of geometry of numbers, a ring of algebraic integers is viewed as a lattice in the $m$-dimensional Euclidean space $\mathbb{R}^{m}$ (see [26, §1.5]).
We let $\sigma_{1},\cdots,\sigma_{s}:K\hookrightarrow\mathbb{R}$ be the $s$ real embeddings and $\sigma_{s+1},\overline{\sigma_{s+1}},\cdots,\sigma_{s+t},\overline{\sigma_{s+t}}:K\hookrightarrow\mathbb{C}$ the $t$ pairs of complex conjugate embeddings of $K$. For each integer $a\in O_{K}$, let $x_{i}=\sigma_{i}(a)$ for $1\leq i\leq s$ when $s\geq 1$, and $y_{j}+\boldsymbol{i}z_{j}=\sigma_{s+j}(a)$ for $1\leq j\leq t$ when $t\geq 1$, where $\boldsymbol{i}$ denotes the imaginary unit in the field $\mathbb{C}$ of complex numbers. Then, $a\in O_{K}$ can be mapped to a point $x=(x_{1},\dotsc,x_{m})=(x_{1},\dotsc,x_{s},y_{1},z_{1},\dotsc,y_{t},z_{t})$ in $\mathbb{R}^{m}$, denoted by $x=\Lambda(a):O_{K}\to\mathbb{R}^{m}$. We have $y_{j}=x_{s+2j-1}$ and $z_{j}=x_{s+2j}$ for $1\leq j\leq t$.
The image of $O_{K}$ under the map $\Lambda:O_{K}\to\mathbb{R}^{m}$ is denoted by $\Lambda_{K}$. Then, $\Lambda_{K}$ is a lattice in $\mathbb{R}^{m}$. It means that $\Lambda_{K}$ is an additive subgroup of $(\mathbb{R}^{m},+)$ and it is discrete in the sense that for any bounded subset $U\subset\mathbb{R}^{m}$, the intersection $U\cap\Lambda_{K}$ is a finite subset of $\mathbb{R}^{m}$. Let $T$ be a fundamental domain of the lattice $\Lambda_{K}$. Then, its volume with Lebesgue measure is given by $\mathrm{vol}(T)=2^{-t}\sqrt{|\Delta_{K}|}$ (see, e.g., [39, Theorem 9.4]).
For a finite set $\Omega$, its cardinality is denoted by $|\Omega|$. The following lemma is an averaging argument in geometry of numbers [19], [38].
Lemma 1
(Lenstra–Siegel) If $U$ is a nonempty bounded subset of $\mathbb{R}^{m}$ which is Lebesgue measurable with volume denoted by $\mathrm{vol}(U)$, then there exists $\tau\in T$ such that
$|(\tau+U)\cap\Lambda_{K}|\geq\frac{\mathrm{vol}(U)}{\mathrm{vol}(T)}.$
Proof: Since $\widetilde{T}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-T$ is also a fundamental domain of the lattice $\Lambda_{K}$, by [39, Lemma 6.2] we know that each point $z\in\mathbb{R}^{m}$ lies in exactly one of the cosets $\tau+\Lambda_{K}$ where $\tau\in\widetilde{T}$.
Let $\chi_{U}(z)$ for $z\in\mathbb{R}^{m}$ be the characteristic function of the Lebesgue measurable subset $U\subseteq\mathbb{R}^{m}$ defined as $1$ if $z\in U$ and 00 otherwise.
We evaluate the following integral
[…]
$=$ $\int_{\tau\in\widetilde{T}}\sum_{y\in\Lambda_{K}}\chi_{U}(\tau+y)d\tau$
[…]
$=$ $\mathrm{vol}(U).$
One can choose any $\tau\in\widetilde{T}=-T$, i.e., $-\tau\in T$ such that the cardinality $|(-\tau+U)\cap\Lambda_{K}|$ is maximal, which is at least $\mathrm{vol}(U)/\mathrm{vol}(T)$. The proof of this lemma is complete.  $\Box$
In the following, we introduce a variant of Lenstra’s construction of codes from geometry of numbers [19]. We assume that $q$ and $r$ are two arbitrarily given positive integers satisfying $2\leq r\leq q$.
We choose $U$ in Lemma 1 as the set of $x\in\mathbb{R}^{m}$ with
$0<x_{i}<\rho\stackrel{{\scriptstyle\mathrm{def}}}{{=}}2^{-t/m}r^{G/m}$
for $1\leq i\leq m$, denoted by $U_{G}$, where $G\in\mathbb{N}$ is a parameter. It is clear that $U_{G}$ is an open subset of $\mathbb{R}^{m}$ with $\mathrm{vol}(U_{G})=\rho^{m}=2^{-t}r^{G}.$ By Lemma 1, there exists $\tau\in T$ such that
$|(\tau+U_{G})\cap\Lambda_{K}|\geq\frac{r^{G}}{\sqrt{|\Delta_{K}|}}.$ (9)
We define the following finite subset of $O_{K}$
$\Omega_{G}=\{a\in O_{K}\mid\Lambda(a)\in(\tau+U_{G})\cap\Lambda_{K}\}$
whose cardinality is equal to $|(\tau+U_{G})\cap\Lambda_{K}|$. By (9), we have $|\Omega_{G}|\geq r^{G}/\sqrt{|\Delta_{K}|}$.
Let $P$ be a nonzero prime ideal of $O_{K}$. Then, the quotient ring $O_{K}/P$ is a finite field, which is called the residue class field of $O_{K}$ modulo $P$. The norm of $P$ is defined by $\mathfrak{N}(P)=|O_{K}/P|$. We define the number
$N_{r,q}(K)=|\{P\subseteq O_{K}\mid r\leq\mathfrak{N}(P)\leq q\}|$ (10)
where $P$ runs through the nonzero prime ideals of $O_{K}$. Note that $N_{r,q}(K)$ is the cardinality of the given subset which has finitely many elements.
In the case that $n=N_{r,q}(K)\geq 1$, we let $P_{1},\cdots,P_{n}$ denote the distinct nonzero prime ideals of $O_{K}$ whose norms are between $r$ and $q$. It is ready to define the following map $\psi:\Omega_{G}\to\mathbb{Z}_{q}^{n}$
$\psi(a)=(a(P_{1}),\dotsc,a(P_{n})),\qquad\mbox{for all }a\in\Omega_{G}$
where $a(P_{i})=a+P_{i}\in O_{K}/P_{i}$ is the residue class of $a\in O_{K}$ modulo $P_{i}$ for $1\leq i\leq n$. By choosing an arbitrary injective map $O_{K}/P_{i}\to\mathbb{Z}_{q}$ for each $i$ with $1\leq i\leq n$, we understand that $\psi(a)\in\mathbb{Z}_{q}^{n}$.
The image of the map $\psi:\Omega_{G}\to\mathbb{Z}_{q}^{n}$, denoted by ${\cal C}(P_{1},\cdots,P_{n};G)$, is a subset of $\mathbb{Z}_{q}^{n}$, which is called a $q$-ary code from geometry of numbers or a $q$-ary Lenstra code. It has some properties as described below.
Proposition 1
Let $K$ be an algebraic number field and $r,q\in\mathbb{N}$ satisfying $2\leq r\leq q$. Assume $n=N_{r,q}(K)\geq 1$. If $r^{n}\geq\sqrt{|\Delta_{K}|}$, then for any $G\in\mathbb{N}$ with $r^{G}\geq\sqrt{|\Delta_{K}|}$ and $G\leq n$, the Lenstra code ${\cal C}(P_{1},\cdots,P_{n};G)$ is a $q$-ary $(n,M,d)$ code with $M\geq r^{G}/\sqrt{|\Delta_{K}|}$ and $d\geq n+1-G$.
Proof: Let $a,b\in\Omega_{G}\subseteq O_{K}$ with $a\neq b$, whose corresponding codewords are $c_{a}=\psi(a)\in\mathbb{Z}_{q}^{n}$ and $c_{b}=\psi(b)\in\mathbb{Z}_{q}^{n}$, respectively.
Let $w=d_{\mathrm{H}}(c_{a},c_{b})$ be the Hamming distance between $c_{a}$ and $c_{b}$. Then, $w=n-\ell$, where $\ell$ is the number of $P_{i}$’s, $1\leq i\leq n$, such that $a(P_{i})=a+P_{i}=b(P_{i})=b+P_{i}$. Without loss of generality, we assume $P_{1},\cdots,P_{\ell}$ being those nonzero prime ideals. Hence, $a-b\in P_{i}$ for $1\leq i\leq\ell$.
Now, each $P_{i}$, $1\leq i\leq\ell$, is a prime factor of the prime ideal factorization in $O_{K}$ of the nonzero principal ideal $(a-b)O_{K}\subseteq O_{K}$. As a consequence, the norm of $(a-b)O_{K}$, defined as the cardinality of the quotient ring $O_{K}/((a-b)O_{K})$, is at least $\prod_{i=1}^{\ell}\mathfrak{N}(P_{i}).$
Note that the norm of the nonzero principal ideal $(a-b)O_{K}$ is equal to the absolute value of the field norm $\mathrm{N}_{K/\mathbb{Q}}(a-b)$ [39, Corollary 5.10]. Thus, we have
$|\mathrm{N}_{K/\mathbb{Q}}(a-b)|\geq\prod_{i=1}^{\ell}\mathfrak{N}(P_{i})\geq r^{\ell}.$ (11)
Let $x=\Lambda(a)$ with $x=(x_{1},\dotsc,x_{m})=(x_{1},\dotsc,x_{s},y_{1},z_{1},\dotsc,y_{t},z_{t})$ and $\hat{x}=\Lambda(b)$ with $\hat{x}=(\hat{x}_{1},\dotsc,\hat{x}_{m})=(\hat{x}_{1},\dotsc,\hat{x}_{s},\hat{y}_{1},\hat{z}_{1},\dotsc,\hat{y}_{t},\hat{z}_{t})$ under the map $\Lambda:O_{K}\to\mathbb{R}^{m}$.
Since $a,b\in\Omega_{G}$, we have $x,\hat{x}\in(\tau+U_{G})\cap\Lambda_{K}$, where $\tau=(\tau_{1},\dotsc,\tau_{m})\in T$ is given in (9). Therefore, $x-\tau,\hat{x}-\tau\in U_{G}$, namely, $0<x_{i}-\tau_{i}<\rho$ and $0<\hat{x}_{i}-\tau_{i}<\rho$ for $1\leq i\leq m$. Hence, $|x_{i}-\hat{x}_{i}|=|(x_{i}-\tau_{i})-(\hat{x}_{i}-\tau_{i})|<\rho$ for $1\leq i\leq m$ and $|y_{j}-\hat{y}_{j}|^{2}+|z_{j}-\hat{z}_{j}|^{2}=|x_{s+2j-1}-\hat{x}_{s+2j-1}|^{2}+|x_{s+2j}-\hat{x}_{s+2j}|^{2}<2\rho^{2}$ for $1\leq j\leq t$.
We estimate the absolute value of the norm $\mathrm{N}_{K/\mathbb{Q}}(a-b)$ as follows
$|\mathrm{N}_{K/\mathbb{Q}}(a-b)|$ $=$ $\prod_{i=1}^{s}|\sigma_{i}(a-b)|\times\prod_{j=1}^{t}|\sigma_{s+j}(a-b)|^{2}$ (12)
$=$ $\prod_{i=1}^{s}|\sigma_{i}(a)-\sigma_{i}(b)|\times\prod_{j=1}^{t}|\sigma_{s+j}(a)-\sigma_{s+j}(b)|^{2}$
$=$ $\prod_{i=1}^{s}|x_{i}-\hat{x}_{i}|\times\prod_{j=1}^{t}|(y_{j}+\boldsymbol{i}z_{j})-(\hat{y}_{j}+\boldsymbol{i}\hat{z}_{j})|^{2}$
$=$ $\prod_{i=1}^{s}|x_{i}-\hat{x}_{i}|\times\prod_{j=1}^{t}[|y_{j}-\hat{y}_{j}|^{2}+|z_{j}-\hat{z}_{j}|^{2}]$
$<$ $\rho^{s}\times(2\rho^{2})^{t}=r^{G}.$
Then, it follows from (11) and (12) that $r^{G}>r^{\ell}$, i.e., $G\geq\ell+1$. Hence, $w=n-\ell\geq n+1-G$. Since $G\leq n$, we have $d_{\mathrm{H}}(c_{a},c_{b})=w\geq 1$, which implies that $c_{a}\neq c_{b}$. Thus, the map $\psi:\Omega_{G}\to\mathbb{Z}_{q}^{n}$ is injective.
Consequently, the code ${\cal C}(P_{1},\cdots,P_{n};G)$, defined as the image of the map $\psi:\Omega_{G}\to\mathbb{Z}_{q}^{n}$, has the size of $M=|{\cal C}(P_{1},\cdots,P_{n};G)|=|\Omega_{G}|\geq r^{G}/\sqrt{|\Delta_{K}|}$ and the minimum Hamming distance of $d\geq n+1-G$. Therefore, the proof is complete.  $\Box$
The code ${\cal C}(P_{1},\cdots,P_{n};G)$ in Proposition 1 is actually a punctured version [20, p. 28] of the code from geometry of numbers originally introduced by Lenstra [19] in 1986. The length of the code constructed in [19] is equal to
$\hat{N}_{r,q}(K)=s+t+|\{P\subseteq O_{K}\mid r\leq\mathfrak{N}(P)^{k(P)}\leq q\;\mbox{for some }k(P)\in\mathbb{N}\}|$ (13)
where $P$ runs through the nonzero prime ideals of $O_{K}$. It is clear that $\hat{N}_{r,q}(K)\geq s+t+N_{r,q}(K)\geq N_{r,q}(K)$. The Lenstra code in [19] has involved all real and complex embeddings of the number field $K$ and possibly more nonzero prime ideals $P\subseteq O_{K}$ in the construction. The code ${\cal C}(P_{1},\cdots,P_{n};G)$ in Proposition 1 is constructed via puncturing the original Lenstra code by keeping only coordinates of each codeword at prime ideals with their norms between $r$ and $q$. The punctured Lenstra code is employed because in this case theory of codes from geometry of numbers is simpler, and it is adequate for our purpose of establishing Theorem 1 and Theorem 8.
It is noticed that Lenstra [19] only gave a complete proof of the assertion regarding properties of $(n,M,d)$ for codes from geometry of numbers in the case of $K$ with $s=[K:\mathbb{Q}]$ and $t=0$, namely, $K$ being a totally real field. Nonetheless, we have given a proof of Proposition 1 as to properties of the simpler punctured Lenstra codes for an arbitrary number field $K$ with $t=0$ or $t>0$. Both cases of $t=0$ and $t>0$ will be needed for the code constructions in the paper.
It is known that the discriminant $\Delta_{K}$ of a number field $K$ is always a nonzero rational integer. We introduce the number
$N_{r,q}(\Delta)=\max_{K:|\Delta_{K}|=\Delta}N_{r,q}(K)$
for any given $\Delta\in\mathbb{N}$, which is well-defined by the Hermite theorem [26, Theorem 3.2.16] asserting that there are only finitely many algebraic number fields $K$, up to isomorphism, with $|\Delta_{K}|\leq\Delta$. Then, for $r,q\in\mathbb{N}$ with $2\leq r\leq q$, we can define the quantity
$A(r,q)=\limsup_{\Delta\rightarrow\infty}\frac{N_{r,q}(\Delta)}{\log\sqrt{\Delta}}.$ (14)
It is clear that for $r_{1},r_{2},q\in\mathbb{N}$ with $2\leq r_{1}\leq r_{2}\leq q$, we have $A(r_{1},q)\geq A(r_{2},q)$. In general, $A(r,q)$ has the following upper bound.
Lemma 2
For all $r,q\in\mathbb{N}$ with $2\leq r\leq q$, we have
$0\leq A(r,q)<7\frac{q}{\log q}.$
Proof: It is obvious that $A(r,q)\geq 0$. We can assume that $A(r,q)>0$.
Let $K$ be an algebraic number field. For each prime ideal $P$ in the subset given in (10) defining the number $N_{r,q}(K)$, it lies over some prime number $p$ in the sense that $P\cap\mathbb{Z}=p\mathbb{Z}$. Then, $\mathfrak{N}(P)=p^{f}$, where $f\geq 1$ is the inertia degree of $P$ over $p\mathbb{Z}$. Since $\mathfrak{N}(P)\leq q$, we have $p\leq q$. On the other hand, by the fundamental identity in extension of algebraic number fields [26, Proposition 1.8.2], there are at most $[K:\mathbb{Q}]$ prime ideals of $O_{K}$ lying over $p\mathbb{Z}$. Therefore, we obtain
$N_{r,q}(K)\leq[K:\mathbb{Q}]\pi(q)$ (15)
where $\pi(q)$ is the prime-counting function.
We need Minkowski’s bound on the discriminant [26, Proposition 3.2.14]
$\sqrt{|\Delta_{K}|}\geq\frac{m^{m}}{m!}\left(\frac{\pi}{4}\right)^{m/2}$ (16)
where $m=[K:\mathbb{Q}]$.
Let $K_{i}$, $i\in\mathbb{N}$, be a sequence of number fields such that $\Delta_{i}=|\Delta_{K_{i}}|>1$ and $N_{r,q}(\Delta_{i})=N_{r,q}(K_{i})$, which satisfy
$\lim_{i\to\infty}\Delta_{i}=\infty\mbox{ and }\lim_{i\to\infty}\frac{N_{r,q}(K_{i})}{\log\sqrt{\Delta_{i}}}=A(r,q)>0.$
This implies $N_{r,q}(K_{i})\to\infty$, and hence, from (15), that $m_{i}=[K_{i}:\mathbb{Q}]\to\infty$ as $i\to\infty$. Therefore, by using Minkowski’s bound (16) and Stirling’s formula for factorials, we get
$A(r,q)\leq\frac{1}{1-\log\frac{2}{\sqrt{\pi}}}\pi(q)=1.1373\cdots\pi(q).$ (17)
Since $\pi(q)<6\frac{q}{\log q}$ for all $q\geq 2$ [1, Theorem 4.6], the proof of this lemma is finished.  $\Box$
Lemma 3
For all positive integers $q\geq 2$, we have
$0\leq A(q,q)\leq\frac{1}{1-\log\frac{2}{\sqrt{\pi}}}=1.1373\cdots.$
Proof: We employ the above proof of Lemma 2. In the specific case of $r=q$, there is either no any or only one prime number $p$ such that $\mathfrak{N}(P)=p^{f}=q$ for the given $q$. Therefore, we obtain
$N_{r,q}(K)\leq[K:\mathbb{Q}]$
instead of (15). Then, the upper bound in this lemma follows, instead of (17).  $\Box$
With the quantity $A(r,q)$ defined in (14), the Lenstra code with properties as detailed in Proposition 1 can provide the following lower bound for the rate function $R_{q}(\delta)$.
Proposition 2
For all $r,q\in\mathbb{N}$ with $2\leq r\leq q$, we have
$R_{q}(\delta)\geq(1-\delta)\frac{\log r}{\log q}-\frac{1}{A(r,q)\log q}$ (18)
for $10<\delta<$ 1.
Proof: We can assume that
$0<\delta<1-\frac{1}{A(r,q)\log r}$
for otherwise the right-hand side of (18) is less than or equal to zero.
Let $K_{i}$, $i\in\mathbb{N}$, be a sequence of number fields such that $\Delta_{i}=|\Delta_{K_{i}}|>1$ and $N_{r,q}(\Delta_{i})=N_{r,q}(K_{i})$ satisfying
$\lim_{i\to\infty}\Delta_{i}=\infty\mbox{ and }\lim_{i\to\infty}\frac{N_{r,q}(K_{i})}{\log\sqrt{\Delta_{i}}}=A(r,q)>\frac{1}{\log r}.$ (19)
In what follows, we consider for $i\in\mathbb{N}$ sufficiently large.
Let $n_{i}=N_{r,q}(K_{i})$. Then, from (19), $n_{i}\to\infty$ as $i\to\infty$. Moreover, $r^{n_{i}}>\sqrt{|\Delta_{K_{i}}|}$.
We choose $G_{i}=\lceil(1-\delta)n_{i}\rceil\in\mathbb{N}$. Then, $G_{i}<n_{i}$. It is clear that
$\lim_{i\to\infty}\frac{G_{i}}{\log\sqrt{|\Delta_{K_{i}}|}}=\lim_{i\to\infty}\frac{G_{i}}{n_{i}}\frac{N_{r,q}(K_{i})}{\log\sqrt{\Delta_{i}}}=(1-\delta)A(r,q)>\frac{1}{\log r}.$
Therefore, $r^{G_{i}}>\sqrt{|\Delta_{K_{i}}|}$ for $i\in\mathbb{N}$ sufficiently large.
By Proposition 1, the Lenstra code ${\cal C}(P_{1},\cdots,P_{n_{i}};G_{i})$ is a $q$-ary $(n_{i},M_{i},d_{i})$ code with
$M_{i}\geq r^{G_{i}}/\sqrt{|\Delta_{K_{i}}|}$
and
$d_{i}\geq n_{i}+1-G_{i}\geq\delta n_{i}$
for $i\in\mathbb{N}$ sufficiently large. Therefore, we have $\alpha_{q}(n_{i},\delta n_{i})\geq r^{G_{i}}/\sqrt{|\Delta_{K_{i}}|}$, and hence
$R_{q}(\delta)$ $=$ $\limsup_{n\rightarrow\infty}\frac{1}{n}\log_{q}\alpha_{q}(n,\delta n)$
$\geq$ $\limsup_{i\to\infty}\frac{1}{n_{i}}\log_{q}\alpha_{q}(n_{i},\delta n_{i})$
$\geq$ $\lim_{i\to\infty}\frac{1}{\log q}\left[\frac{G_{i}}{n_{i}}\log r-\frac{\log\sqrt{|\Delta_{K_{i}}|}}{n_{i}}\right]$
$=$ $(1-\delta)\frac{\log r}{\log q}-\frac{1}{A(r,q)\log q}.$
The proof of this proposition is complete.  $\Box$
By choosing $\delta=1-q^{-1}$ for which $R_{q}(\delta)=0$, Proposition 2 can yield the following upper bound of $A(r,q)$, which depends on both $r$ and $q$.
Corollary 1
For all $r,q\in\mathbb{N}$ with $2\leq r\leq q$, we have
$0\leq A(r,q)\leq\frac{q}{\log r}.$
It is interesting that this new upper bound of $A(r,q)$ cannot be subsumed by (17), since it is better than (17) in the case of $r=\lfloor q^{0.99}\rfloor$ and $q$ being sufficiently large, as seen by the prime number theorem [1, p. 74].
From (18) in Proposition 2, we observe that the Lenstra code would provide a lower bound on the rate function $R_{q}(\delta)$, if a lower bound on $A(r,q)$, which is defined by (14), were found.
We are going to make use of infinite towers of Hilbert class fields in class field theory for algebraic number fields to achieve a lower bound for the quantity $A(r,q)$.
3 Hilbert class field tower and its concatenation with Lenstra codes
It was an open question in algebraic number theory, going back to Furtwängler and Hilbert, whether every number field can be embedded in a number field with class number being $1$ (i.e., equal to its Hilbert class field). In 1964, Golod and Shafarevich [11] provided a negative answer to this question by showing that the Hilbert class field tower of a number field $K$ may be infinite. An explicit example of such number field given in [11, p. 270] is a quadratic field $K=\mathbb{Q}(\sqrt{-3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 19})=\mathbb{Q}(\sqrt{-4849845})$.
Serre [33], [34] firstly presented a surprising application of an infinite Hilbert class field tower, with prescribed places splitting completely, for algebraic function fields of curves over finite fields to obtain a general lower bound on Ihara’s number $A(q)$ defined by (1) (see also [27] and [32] for expositions).
In this section, we follow the idea of Serre to show that one can apply an infinite tower of Hilbert class fields, with prescribed places splitting completely, for algebraic number fields to obtain a lower bound for the quantity $A(r,q)$ defined by (14). Then, a lower bound for $A(r,q)$ attainable from infinite Hilbert class field towers can yield, through Lenstra codes, a lower bound for the rate function $R_{q}(\delta)$.
3.1 Hilbert class field tower
We give some preliminaries about infinite towers of Hilbert class fields of an algebraic number field in class field theory.
A place (or prime) $P$ of an algebraic number field $K$ is a class of equivalent valuations, or a normalized valuation, of $K$. The archimedean equivalence classes are called infinite places, and the non-archimedean ones called finite places. Let $\mathbb{P}_{K}$ denote the set of places of $K$, $S_{\infty}$ the set of infinite, or archimedean, places of $K$, and $S_{f}$ the set of finite, or non-archimedean, places of $K$. Thus, $\mathbb{P}_{K}=S_{\infty}\cup S_{f}$. An infinite place $P\in\mathbb{P}_{K}$ is referred to by the notation $P\mid\infty$ and a finite one by $P\nmid\infty$. It is known that the infinite places $P\mid\infty$ of $K$ are obtained from the real and complex embeddings of $K$. There are $s$ real places associated with the $s$ real embeddings of $K$ and $t$ complex places with the $t$ pairs of complex conjugate embeddings of $K$. Thus, the number of infinite places of $K$ is given by $|S_{\infty}|=s+t$. The finite places $P\nmid\infty$ of $K$ can stand for the nonzero prime ideals $P$ of $O_{K}$ by the normalized $P$-adic exponential valuation (see, e.g., [26]).
Let $I_{K}$ denote the multiplicative group of nonzero fractional ideals of $K$, and $P_{K}$ its subgroup of nonzero principal fractional ideals of $K$. The quotient group $I_{K}/P_{K}$, denoted by $Cl_{K}$, is called the ideal class group of $K$, which is a finite abelian group. The order of the ideal class group $Cl_{K}$ is called the class number of $K$, denoted by $h_{K}$. Let $U_{K}$ be the multiplicative group of units of the ring of integers $O_{K}$, which is a subgroup of the multiplicative group $K^{*}=K\setminus\{0\}$. It is noteworthy to remark that for the two multiplicative groups of $K^{*}$ and $I_{K}$, there is a group homomorphism $f:K^{*}\to I_{K}$ sending $a\in K^{*}$ to the principal fractional ideal $aO_{K}\in I_{K}$. Then, $U_{K}$ is the kernel of $f$ denoted by $\mathrm{Ker}f$, $P_{K}$ the image of $f$ by $\mathrm{Im}f$, and $Cl_{K}$ the cokernel of $f$ by $\mathrm{Coker}f$.
A main result of class field theory is that, for an algebraic number field $K$ of finite degree, there exists a maximal unramified abelian extension of $K$, denoted by $H_{K}$, in which the real places of $K$ stay real. The number field $H_{K}$ is called the Hilbert class field of $K$ (see [26, p. 399]). The most important property of $H_{K}$ is that the Galois group of the extension $H_{K}/K$ is isomorphic to the ideal class group $Cl_{K}$, i.e., $\mathrm{Gal}(H_{K}/K)\cong Cl_{K}$. There is no finite place $P\nmid\infty$ of $K$ ramified in $H_{K}$. Moreover, if $\sigma:K\hookrightarrow\mathbb{R}$ is a real embedding of $K$, then all extensions of $\sigma$ to $H_{K}$ remain real. It is clear that $H_{K}=K$ if and only if $h_{K}=1$ (i.e., $O_{K}$ is a principal ideal domain).
We are ready to define a sequence of field extensions as follows
$K=K_{1}\subseteq K_{2}\subseteq K_{3}\subseteq\cdots$
where $K_{i+1}=H_{K_{i}}$, the Hilbert class field of $K_{i}$, for $i\in\mathbb{N}$. This sequence of fields is called the Hilbert class field tower of $K$. We say that $K$ has an infinite Hilbert class field tower if $[K_{i}:K]\to\infty$ as $i\to\infty$ and finite otherwise.
We need a notion of $p$-rank in group theory [27]. If $G$ is a finitely generated abelian group (written additively) and $p$ is a prime, the $p$-rank of $G$, denoted by $d_{p}G$, is defined as the dimension of the $\mathbb{F}_{p}$-vector space $G/pG$. If $G$ is a finite abelian group, then the $p$-rank $d_{p}G$ is equal to the number of summands in the decomposition of the $p$-Sylow subgroup of $G$ as a direct sum of cyclic subgroups. In general, a finitely generated abelian group $G$ can be decomposed as a direct sum of $G=G_{t}\oplus G_{f}$, where $G_{t}$ is the torsion subgroup of finite order of $G$, and $G_{f}$ is a free abelian subgroup of $G$. Then, the $p$-rank $d_{p}G$ is the sum of the $p$-rank $d_{p}G_{t}$ plus the rank of the free abelian group $G_{f}$.
The Golod–Shafarevich condition for the infinite Hilbert class field tower of $K$ (see Roquette [29, p. 235]) asserts that if, for some prime $p$,
$d_{p}Cl_{K}\geq 2+2\sqrt{d_{p}U_{K}+1}$ (20)
then $K$ has an infinite Hilbert class field tower.
By Dirichlet’s unit theorem [39, p. 300], $U_{K}$ is a finitely generated abelian group and it follows that
$d_{p}U_{K}=s+t-1+\delta_{p}(K)=|S_{\infty}|-1+\delta_{p}(K)$ (21)
where $\delta_{p}(K)$ equals $1$ if $K$ contains a primitive $p$th root of unity and 00 otherwise.
It is indicated by (20) that the ideal class group $Cl_{K}$ and the group of units $U_{K}$ of $K$ can jointly determine a situation to guarantee infiniteness of the Hilbert class field tower of $K$. By virtue of the above mentioned homomorphism $f:K^{*}\to I_{K}$, the Golod–Shafarevich condition (20) can be rewritten as
$d_{p}\mathrm{Coker}f\geq 2+2\sqrt{d_{p}\mathrm{Ker}f+1}.$
The above theory of infinite Hilbert class field towers can be generalized to including a requirement of some prescribed finite places splitting completely in the class field. The generalized theory can then be connected with Lenstra codes to provide a new lower bound for the rate function $R_{q}(\delta)$. We need to use the concepts of adèle and idèle in the generalized theory. The two groups of $Cl_{K}$ and $U_{K}$ in the Golod–Shafarevich condition (20) will be replaced by a generalized version of theirs.
3.2 $S_{c}$-Hilbert class field tower
Let $K$ be an algebraic number field and $S_{c}$ a finite subset of finite places of $K$, i.e., $S_{c}\subseteq S_{f}\subseteq\mathbb{P}_{K}$ with $|S_{c}|<\infty$. In the following, we present a route for theory of $S_{c}$-Hilbert class field of $K$ (see Tate [41], Cassels [5], and Neukirch [26, Chapter VI]).
For each $P\in\mathbb{P}_{K}$, let $K_{P}$ denote the completion of the global field $K$ and $K_{P}^{*}=K_{P}\setminus\{0\}$. If $P\mid\infty$ is a real or complex place, then $K_{P}=\mathbb{R}$ or $\mathbb{C}$, respectively. If $P\nmid\infty$, let $v_{P}$ be the normalized discrete valuation of $K_{P}$ with $v_{P}(K_{P}^{*})=\mathbb{Z}$. For each $P\nmid\infty$, the valuation ring of $K_{P}$ is denoted by $O_{P}=\{a\in K_{P}\mid v_{P}(a)\geq 0\}$ and the group of units of $O_{P}$ by $U_{P}=\{a\in K_{P}\mid v_{P}(a)=0\}$.
An adèle of $K$ is an element $\alpha=(\alpha_{P})$ of the direct product $\prod_{P\in\mathbb{P}_{K}}K_{P}$ such that the $P$-component $\alpha_{P}\in O_{P}$ for all but finitely many places including infinite ones. The set of adèles of $K$, denoted by $\mathbb{A}_{K}$, is a ring under addition and multiplication defined componentwise. The number field $K$ can be embedded into $\mathbb{A}_{K}$ by the diagonal embedding $K\hookrightarrow\mathbb{A}_{K}$ sending $a\in K$ to $\alpha=(\alpha_{P})\in\mathbb{A}_{K}$ with $\alpha_{P}=a$ for all $P\in\mathbb{P}_{K}$.
The idèle group of $K$, denoted by $J_{K}$, is the multiplicative group of units of $\mathbb{A}_{K}$. An idèle of $K$ is an element $\alpha=(\alpha_{P})$ of the direct product $\prod_{P\in\mathbb{P}_{K}}K_{P}^{*}$ such that the $P$-component $\alpha_{P}\in U_{P}$ for all but finitely many places including infinite ones. The multiplicative group $K^{*}$ can be embedded into the abelian group $J_{K}$ by the diagonal embedding $K^{*}\hookrightarrow J_{K}$ sending $a\in K^{*}$ to $\alpha=(\alpha_{P})\in J_{K}$ with $\alpha_{P}=a$ for all $P\in\mathbb{P}_{K}$. Each element of $K^{*}$, which is viewed as a subgroup of $J_{K}$, is called a principal idèle of $K$. The quotient group $C_{K}=J_{K}/K^{*}$ is called the idèle class group of $K$.
Before we move forward, it is mentioned that the following three group isomorphism theorems [13, p. 44] are useful in our development of theory.
Let $G$ be an abelian group (written multiplicatively). Then, each subgroup of $G$ is a normal subgroup. Let $N$ and $K$ be subgroups of the abelian group $G$. We denote by $NK$ the subset $\{ab\mid a\in N\mbox{ and }b\in K\}\subseteq G$. Then, $NK$ is a subgroup of $G$ and it includes $N$ and $K$ as its subgroups. Let $H$ be a group. The first group isomorphism theorem asserts that if $\phi:G\to H$ is a group homomorphism, then $G/\mathrm{Ker}\phi\cong\mathrm{Im}\phi$. The second group isomorphism theorem claims that $NK/N\cong K/(N\cap K)$. The third group isomorphism theorem says that if $N$ is a subgroup of $K$, then $K/N$ is a normal subgroup of $G/N$ and $(G/N)/(K/N)\cong G/K$.
Note that each finite place $P\in S_{f}$ can be identified by a nonzero prime ideal $P$ of $O_{K}$. There is a surjective homomorphism $\pi_{f}:J_{K}\to I_{K}$ from the idèle group $J_{K}$ to the fractional ideal group $I_{K}$ defined by, for each $\alpha=(\alpha_{P})\in J_{K}$,
$\pi_{f}(\alpha)=\prod_{P\nmid\infty}P^{v_{P}(\alpha_{P})}\in I_{K}.$ (22)
The kernel of the homomorphism $\pi_{f}:J_{K}\to I_{K}$ is given by
$J_{K}^{S_{\infty}}=\prod_{P\mid\infty}K_{P}^{*}\times\prod_{P\nmid\infty}U_{P}.$
For a given nonempty finite subset $S_{c}\subseteq S_{f}$, let $I_{K}^{S_{c}}$ be the subgroup of $I_{K}$ which is finitely generated by the prime ideals in $S_{c}$. If $S_{c}=\emptyset$, then $I_{K}^{S_{c}}$ is the trivial group, which has only one element being $O_{K}$. Now, the composite of $\pi_{f}:J_{K}\to I_{K}$ and the canonical projection $\pi_{S_{c}}:I_{K}\to I_{K}/I_{K}^{S_{c}}$ can define a group homomorphism from $J_{K}$ to $I_{K}/I_{K}^{S_{c}}$ as follows
$\pi_{f,S_{c}}:J_{K}\to I_{K}\to I_{K}/I_{K}^{S_{c}}.$ (23)
Let $S=S_{\infty}\cup S_{c}$. The kernel of the above surjective homomorphism $\pi_{f,S_{c}}:J_{K}\to I_{K}/I_{K}^{S_{c}}$ is given by
$J_{K}^{S}=\prod_{P\in S}K_{P}^{*}\times\prod_{P\notin S}U_{P}.$ (24)
By applying the first group isomorphism theorem to the homomorphism $\pi_{f,S_{c}}$, we obtain
$J_{K}/J_{K}^{S}=J_{K}/\mathrm{Ker}\pi_{f,S_{c}}\cong\mathrm{Im}\pi_{f,S_{c}}=I_{K}/I_{K}^{S_{c}}.$ (25)
By the unique prime ideal factorization theorem [26, Theorem 1.3.3] and its relation with exponential valuation [26, p. 69], the group homomorphism $f:K^{*}\to I_{K}$ sending $a\in K^{*}$ to $aO_{K}\in I_{K}$ is the same as the composite of the diagonal embedding $K^{*}\hookrightarrow J_{K}$ and the group homomorphism $\pi_{f}:J_{K}\to I_{K}$ given in (22). Therefore, the image of the following composite homomorphism
$\hat{f}:K^{*}\hookrightarrow J_{K}\to I_{K}\to I_{K}/I_{K}^{S_{c}}$ (26)
is the same as the image of the composite homomorphism of $f:K^{*}\to I_{K}$ and $\pi_{S_{c}}:I_{K}\to I_{K}/I_{K}^{S_{c}}$, namely, $K^{*}\to I_{K}\to I_{K}/I_{K}^{S_{c}}$. Hence, the image of $\hat{f}$ is identical to the image of the following composite homomorphism
$\tilde{f}:P_{K}\hookrightarrow I_{K}\to I_{K}/I_{K}^{S_{c}}$ (27)
where $P_{K}$ is the image of $f:K^{*}\to I_{K}$, and $P_{K}\hookrightarrow I_{K}$ is the canonical injection. That is to say $\mathrm{Im}\hat{f}=\mathrm{Im}\tilde{f}$.
Noting that the homomorphism $\hat{f}:K^{*}\to I_{K}/I_{K}^{S_{c}}$ given in (26) is also a composite of the diagonal embedding $K^{*}\hookrightarrow J_{K}$ and the homomorphism $\pi_{f,S_{c}}:J_{K}\to I_{K}/I_{K}^{S_{c}}$ given in (23), we have $\mathrm{Ker}\hat{f}=K^{*}\cap\mathrm{Ker}\pi_{f,S_{c}}=J_{K}^{S}\cap K^{*}$. It is clear that $\mathrm{Ker}\tilde{f}=P_{K}\cap\mathrm{Ker}\pi_{S_{c}}=I_{K}^{S_{c}}\cap P_{K}$. Then, by using the first group isomorphism theorem, we get
$K^{*}/(J_{K}^{S}\cap K^{*})=K^{*}/\mathrm{Ker}\hat{f}\cong\mathrm{Im}\hat{f}=\mathrm{Im}\tilde{f}\cong P_{K}/\mathrm{Ker}\tilde{f}=P_{K}/(I_{K}^{S_{c}}\cap P_{K}).$ (28)
We define $C_{K}^{S}=J_{K}^{S}K^{*}/K^{*}$. Since $K^{*}\subseteq J_{K}^{S}K^{*}\subseteq J_{K}$, we know that $C_{K}^{S}$ is a normal subgroup of $J_{K}/K^{*}=C_{K}$. Similarly, we let $Cl_{K}^{S_{c}}=I_{K}^{S_{c}}P_{K}/P_{K}$ and, by $P_{K}\subseteq I_{K}^{S_{c}}P_{K}\subseteq I_{K}$, we see that $Cl_{K}^{S_{c}}$ is a normal subgroup of $I_{K}/P_{K}=Cl_{K}$. If $S_{c}=\emptyset$, then $Cl_{K}^{S_{c}}$ is the trivial group. There is a group isomorphism as shown below.
Proposition 3
For any finite subset $S_{c}\subseteq S_{f}$, let $S=S_{\infty}\cup S_{c}$. Then, we have
$C_{K}/C_{K}^{S}\cong Cl_{K}/Cl_{K}^{S_{c}}.$ (29)
Proof: By the second and third group isomorphism theorems, we obtain $C_{K}/C_{K}^{S}=(J_{K}/K^{*})/(J_{K}^{S}K^{*}/K^{*})\cong J_{K}/J_{K}^{S}K^{*}\cong(J_{K}/J_{K}^{S})/(J_{K}^{S}K^{*}/J_{K}^{S})\cong(J_{K}/J_{K}^{S})/(K^{*}/(J_{K}^{S}\cap K^{*}))\cong(I_{K}/I_{K}^{S_{c}})/(P_{K}/(I_{K}^{S_{c}}\cap P_{K}))\cong(I_{K}/I_{K}^{S_{c}})/(I_{K}^{S_{c}}P_{K}/I_{K}^{S_{c}})\cong I_{K}/I_{K}^{S_{c}}P_{K}\cong Cl_{K}/Cl_{K}^{S_{c}},$ where we have used (25) and (28). The proof is finished.  $\Box$
It follows from the finiteness of the ideal class group $Cl_{K}$ and the group isomorphism given in (29) that $C_{K}^{S}$ is a subgroup of $C_{K}$ of finite index. According to [26, Proposition 6.1.8], the subgroup $C_{K}^{S}$ of $C_{K}$ satisfies the topological requirement in the assumption of the existence theorem in the global class field theory (see, [26, Theorem 6.6.1] and [41, p. 172]). Therefore, there exists a unique finite abelian extension of $K$, denoted by $H_{K,S_{c}}$, corresponding to the subgroup $C_{K}^{S}$ of $C_{K}$. The Galois group of the abelian extension $H_{K,S_{c}}/K$ is given by
$\mathrm{Gal}(H_{K,S_{c}}/K)\cong C_{K}/C_{K}^{S}\cong Cl_{K}/Cl_{K}^{S_{c}}.$
The number field $H_{K,S_{c}}$ is called the $S_{c}$-Hilbert class field of $K$. If $S_{c}=\emptyset$, then the subgroup $C_{K}^{S}=C_{K}^{S_{\infty}}=J_{K}^{S_{\infty}}K^{*}/K^{*}$ and $H_{K,S_{c}}=H_{K}$, the Hilbert class field of $K$.
For a finite place $P\nmid\infty$, the multiplicative group $K_{P}^{*}$ can be embedded into $J_{K}$ by the local embedding $K_{P}^{*}\hookrightarrow J_{K}$ sending $a\in K_{P}^{*}$ to $\alpha=(\alpha_{P})\in J_{K}$ with $\alpha_{P}=a$ and $\alpha_{P^{\prime}}=1$ for all $P^{\prime}\in\mathbb{P}_{K}$ with $P^{\prime}\neq P$. Under the local embedding $K_{P}^{*}\hookrightarrow J_{K}$, it follows from (24) and $U_{P}\subseteq K_{P}^{*}$ that $K_{P}^{*}\subseteq J_{K}^{S}K^{*}$ for all $P\in S_{c}\subseteq S$ and $U_{P}\subseteq J_{K}^{S}K^{*}$ for all $P\nmid\infty$ for the subgroup $C_{K}^{S}=J_{K}^{S}K^{*}/K^{*}$ of $C_{K}$.
By using the criteria given in [27, Proposition 2.5.3], which apply for all non-archimedean places of global fields including finite places of algebraic number fields, we see that the $S_{c}$-Hilbert class field $H_{K,S_{c}}$ of $K$ has the following two properties
1) Each finite place $P\nmid\infty$ of $K$ is unramified in $H_{K,S_{c}}$, and
2) Each finite place $P\in S_{c}$ splits completely in $H_{K,S_{c}}$.
It is ready to build a tower of fields as follows
$K=K_{1}^{(S_{1})}\subseteq K_{2}^{(S_{2})}\subseteq K_{3}^{(S_{3})}\subseteq\cdots$
where $S_{1}=S_{c}$ and for $i\in\mathbb{N}$, $K_{i+1}^{(S_{i+1})}=H_{K_{i}^{(S_{i})},S_{i}}$, the $S_{i}$-Hilbert class field of $K_{i}^{(S_{i})}$, and $S_{i+1}$ is the set of all finite places of $K_{i+1}^{(S_{i+1})}$ lying over the places in $S_{i}$. This tower of fields is called the $S_{c}$-Hilbert class field tower of $K$. If $[K_{i}^{(S_{i})}:K]\to\infty$ as $i\to\infty$, we say that $K$ has an infinite $S_{c}$-Hilbert class field tower.
Let $K^{S}=K^{*}\cap J_{K}^{S}=\{a\in K^{*}\mid v_{P}(a)=0\mbox{ for all }P\in S_{f}\setminus S_{c}\}$, each element of which is called an $S$-unit of $K$. If $S_{c}=\emptyset$, i.e., $S=S_{\infty}$, then $K^{S}=K^{S_{\infty}}=K^{*}\cap J_{K}^{S_{\infty}}=U_{K}$, the group of units of $O_{K}$. The generalized Golod–Shafarevich condition has been given by Ihara [17, p. 705], which asserts that if, for some prime $p$,
$d_{p}(Cl_{K}/Cl_{K}^{S_{c}})\geq 2+2\sqrt{d_{p}K^{S}+1}$ (30)
then $K$ has an infinite $S_{c}$-Hilbert class field tower. If $S_{c}=\emptyset$, then $Cl_{K}/Cl_{K}^{S_{c}}\cong Cl_{K}$ and $K^{S}=U_{K}$ and hence that (30) reduces to (20).
It was mentioned in [17, p. 705] that the proof presented in Roquette [29] for the case $S_{c}=\emptyset$ can be generalized immediately to the nonempty case $S_{c}\neq\emptyset$ just by considering the elements of $S_{c}$ as additional “infinite primes”. A proof for an even more general theorem including the generalized Golod–Shafarevich condition (30) has been briefly presented by Maire [21]. The generalized Golod–Shafarevich condition (30) for an infinite class field tower of an algebraic number field is an analogue of the one applied for an infinite class field tower of an algebraic function field over a finite field (see Schoof [32, p. 11], Niederreiter and Xing [27, p. 59], and the monograph by Serre [35]).
For the purpose of applicability, we are going to present a slight generalization of the Golod–Shafarevich condition (20), which will be suitable for application to the code constructions in this paper.
[…]
If $L=H_{K}$, the Hilbert class field of $K$, then the set $P(L/K)$ consists precisely of the nonzero prime ideals of $K$ which are also principal ideals. Since $[H_{K}:K]=h_{K}$, the class number of $K$, the Dirichlet density of the set $P(H_{K}/K)$ is equal to $1/h_{K}$. By the finiteness of $h_{K}$, the set $P(H_{K}/K)$ must have positive density. Therefore, it is guaranteed that one can choose an arbitrarily large nonempty finite subset $S_{c}\subseteq P(H_{K}/K)$.
For any nonempty finite subset $S_{c}\subseteq P(H_{K}/K)$, each of the prime ideals in $S_{c}$ has the form of $aO_{K}$ with $a\in O_{K}$ and $a\neq 0$, which is a principal ideal of $O_{K}$. Hence, $S_{c}\subseteq P_{K}$, the subgroup of nonzero principal fractional ideals of $I_{K}$. Noting that $I_{K}^{S_{c}}$ is the subgroup of $I_{K}$ that is finitely generated by the prime ideals in $S_{c}$, we have $I_{K}^{S_{c}}\subseteq P_{K}$. Thus, $I_{K}^{S_{c}}P_{K}=P_{K}$. Consequently, $Cl_{K}^{S_{c}}=I_{K}^{S_{c}}P_{K}/P_{K}$ is the trivial group.
Therefore, for any finite subset $S_{c}\subseteq P(H_{K}/K)$, it is true that $Cl_{K}/Cl_{K}^{S_{c}}\cong Cl_{K}$ and then the $p$-rank $d_{p}(Cl_{K}/Cl_{K}^{S_{c}})=d_{p}Cl_{K}$ for any prime $p$. By using (30), we have the following criterion.
Proposition 4
Let $S_{c}\subseteq P(H_{K}/K)$ be a finite subset and $S=S_{\infty}\cup S_{c}$. If, for some prime $p$,
$d_{p}(Cl_{K})\geq 2+2\sqrt{d_{p}K^{S}+1}$ (31)
then $K$ has an infinite $S_{c}$-Hilbert class field tower.
The $p$-rank of $K^{S}$, i.e., $d_{p}K^{S}$, in (31) can be calculated as follows. According to a generalized Dirichlet’s unit theorem [26, Proposition 6.1.1], we know that $K^{S}$ is a finitely generated abelian group and that
$d_{p}K^{S}=|S|-1+\delta_{p}(K)=|S_{c}|+|S_{\infty}|-1+\delta_{p}(K)$ (32)
where $\delta_{p}(K)$ equals $1$ if $K$ contains a primitive $p$th root of unity and 00 otherwise.
By employing an infinite tower of $S_{c}$-Hilbert class fields with a constraint on the norms of prime ideals in $S_{c}$, we can obtain a lower bound on the quantity $A(r,q)$. Subsequently, by using Lenstra codes, we can achieve a lower bound on the rate function $R_{q}(\delta)$, as shown below.
Proposition 5
Let $K$ be a number field of discriminant $\Delta_{K}$ with $|\Delta_{K}|>1$. Let $r$ and $q$ be two positive integers satisfying $2\leq r\leq q$ and $S_{c}\subseteq P(H_{K}/K)$ be a nonempty finite subset, each prime ideal in which has norm between $r$ and $q$. If, for some prime $p$,
$d_{p}(Cl_{K})\geq 2+2\sqrt{|S_{c}|+|S_{\infty}|+1}$ (33)
then
$R_{q}(\delta)\geq(1-\delta)\frac{\log r}{\log q}-\frac{\log\sqrt{|\Delta_{K}|}}{|S_{c}|\log q}$ (34)
for $10<\delta<$ 1.
Proof: It is seen from (32) that $d_{p}K^{S}\leq|S_{c}|+|S_{\infty}|$. If, for some prime $p$, the condition (33) is satisfied, then (31) holds. According to Proposition 4, $K$ has an infinite $S_{c}$-Hilbert class field tower.
Let $K=K_{1}^{(S_{1})}\subseteq K_{2}^{(S_{2})}\subseteq\cdots$ be the infinite $S_{c}$-Hilbert class field tower of $K$, where $S_{1}=S_{c}$. Then, $\lim_{i\to\infty}[K_{i}^{(S_{i})}:K]\to\infty$. Moreover, $K_{i+1}^{(S_{i+1})}$ is the $S_{i}$-Hilbert class field of $K_{i}^{(S_{i})}$ for $i\in\mathbb{N}$.
Let $K_{i}=K_{i}^{(S_{i})}$ and $\Delta_{i}=|\Delta_{K_{i}}|$ for $i\in\mathbb{N}$. Then, $[K_{i}:K]\to\infty$ as $i\to\infty$. Since each finite place $P\nmid\infty$ of $K_{i}$ is unramified in $K_{i+1}$, by virtue of an analogue of the Riemann–Hurwitz formula for algebraic number fields [26, Proposition 3.3.13], we have
$\log\sqrt{\Delta_{i+1}}=[K_{i+1}:K_{i}]\log\sqrt{\Delta_{i}}$
for $i\in\mathbb{N}$. Hence,
$\log\sqrt{\Delta_{i}}=[K_{i}:K_{1}]\log\sqrt{|\Delta_{K}|}$ (35)
for $i\in\mathbb{N}$, which implies $\Delta_{i}\to\infty$ as $i\to\infty$.
For each nonzero prime ideal $P_{i}\in S_{i}$, it splits completely in $K_{i+1}$ for $i\in\mathbb{N}$. Therefore, there are $[K_{i+1}:K_{i}]$ nonzero prime ideals in $S_{i+1}$ lying over each $P_{i}\in S_{i}$ for $i\in\mathbb{N}$. Consequently, we have $|S_{i+1}|=[K_{i+1}:K_{i}]|S_{i}|$ for $i\in\mathbb{N}$. Moreover, for $i\in\mathbb{N}$ with $i\geq 2$, each $P_{i}\in S_{i}$ has the same norm as that of the prime ideal $P\in S_{c}$ lying below $P_{i}$. Hence, for $i\in\mathbb{N}$, the norm of each prime ideal $P_{i}\in S_{i}$ is between $r$ and $q$. Then, it follows that
$N_{r,q}(K_{i})\geq|S_{i}|=[K_{i}:K_{1}]|S_{1}|=[K_{i}:K_{1}]|S_{c}|$ (36)
for $i\in\mathbb{N}$.
Therefore, we get
[…]
where the last inequality follows from (35) and (36). It is clear that (34) can be deduced from (18) in Proposition 2 and (37).
The proof of this proposition is complete.  $\Box$
Proposition 5 is the theoretical foundation for the following practical code constructions in this paper. For the application, a lower bound for the $p$-rank of the ideal class group $Cl_{K}$, i.e., $d_{p}Cl_{K}$, in the Golod–Shafarevich condition (33) will be needed.
4 An applicable lower bound for the rate function
Based on theory of Lenstra codes from geometry of numbers given in Section 2 and theory of infinite towers of Hilbert class fields in Section 3, we can obtain an applicable lower bound on the rate function $R_{q}(\delta)$ in the following.
Let $p_{i}\in\mathbb{P}$ be the $i$th prime for $i\in\mathbb{N}$. For an odd prime $p$ and an integer $a\in\mathbb{Z}$, the Legendre symbol $\big{(}\frac{a}{p}\big{)}\in\{-1,0,1\}$ is function of $a$ and $p$ defined as 00 when $p\mid a$ and ±$\pm 1$ depending on whether $a$ is a quadratic residue modulo $p$ when $p\nmid a$.
Theorem 3
Let $q\in\mathbb{N}$ with $q\geq 2$. If there exist $r,\ell,k\in\mathbb{N}$ which satisfy the following three conditions
1) $2\leq r\leq q$,
2) $k+2\leq\frac{1}{4}(\ell-2)^{2}-(\ell-2)$, and
3) There are at least $2k$ primes $p$ satisfying $r\leq p^{2}\leq q$, $p>p_{\ell}$, and $p\equiv 3\pmod{4}$,
then
$R_{q}(\delta)\geq(1-\delta)\frac{\log r}{\log q}-\frac{\log\sqrt{D}}{k\log q}$ (38)
for $10<\delta<$ 1, where $D=4p_{1}\cdots p_{\ell}$.
Proof: We consider the quadratic number field $K=\mathbb{Q}(\sqrt{\Delta_{K}})$, where $\Delta_{K}=D$ or $-D$. It is clear that the discriminant of $K$ is $\Delta_{K}$. If $\Delta_{K}>0$, then $K$ is a real quadratic field. If $\Delta_{K}<0$, then $K$ is an imaginary quadratic field.
Since $\Delta_{K}$ is the discriminant of the quadratic field $K$, it is a fundamental discriminant. Thus, according to [6, Theorem 2.2.15], the Kronecker symbol $\big{(}\frac{\Delta_{K}}{n}\big{)}$, $n\in\mathbb{Z}$, defines a real primitive Dirichlet character modulo $|\Delta_{K}|=D$. We note that if $n=p$, $p$ an odd prime, then the Kronecker symbol $\big{(}\frac{\Delta_{K}}{n}\big{)}$ has the same value as that of the Legendre symbol $\big{(}\frac{\Delta_{K}}{p}\big{)}$. Therefore, if $p$ is an odd prime, then
$\left(\frac{-D}{p}\right)=\left(\frac{-1}{p}\right)\left(\frac{D}{p}\right).$
It is known, by Euler’s criterion, that $\big{(}\frac{-1}{p}\big{)}=-1$ for every prime $p\equiv 3\pmod{4}$. Thus, if $p\equiv 3\pmod{4}$ and $p\nmid D$, then either $\big{(}\frac{D}{p}\big{)}=-1$ or $\big{(}\frac{-D}{p}\big{)}=-1$.
For each of the primes $p\equiv 3\pmod{4}$ in condition 3), we have $p>p_{\ell}$ and thus $p\nmid D$. Therefore, each of the primes $p$ in condition 3) satisfies either $\big{(}\frac{D}{p}\big{)}=-1$ or $\big{(}\frac{-D}{p}\big{)}=-1$. Based on this, condition 3) can be reduced to the following two cases, namely, either
Case 1) There are at least $k$ primes $p\equiv 3\pmod{4}$ satisfying $r\leq p^{2}\leq q$, $p>p_{\ell}$, and $\big{(}\frac{D}{p}\big{)}=-1$, or
Case 2) There are at least $k$ primes $p\equiv 3\pmod{4}$ satisfying $r\leq p^{2}\leq q$, $p>p_{\ell}$, and $\big{(}\frac{-D}{p}\big{)}=-1$.
Now, we can choose the quadratic number field $K=\mathbb{Q}(\sqrt{\Delta_{K}})$, where $\Delta_{K}=D$ in Case 1) and $\Delta_{K}=-D$ in Case 2). It follows from condition 3) that there are at least $k$ primes $p\equiv 3\pmod{4}$ satisfying $r\leq p^{2}\leq q$, $p>p_{\ell}$, and $\big{(}\frac{\Delta_{K}}{p}\big{)}=-1$. By using [6, Proposition 3.4.3], we know that each of these $k$ primes $p$ is inert in $K$, namely, the principal ideal $pO_{K}$ being a nonzero prime ideal of $O_{K}$. Thus, the set $\{pO_{K}\mid r\leq p^{2}\leq q,\;p>p_{\ell},\mbox{ and }\big{(}\frac{\Delta_{K}}{p}\big{)}=-1\}$ has cardinality of at least $k$, and it is a subset of $P(H_{K}/K)$.
Therefore, condition 3) implies that there exists a subset, denoted by $S_{c}$, of the set $\{pO_{K}\mid r\leq p^{2}\leq q,\;p>p_{\ell},\mbox{ and }\big{(}\frac{\Delta_{K}}{p}\big{)}=-1\}$ such that $|S_{c}|=k$ and $S_{c}\subseteq P(H_{K}/K)$.
Note that the norm of each $pO_{K}\in S_{c}$ is given by
$\mathfrak{N}(pO_{K})=|\mathrm{N}_{K/\mathbb{Q}}(p)|=p^{[K:\mathbb{Q}]}=p^{2}.$
Thus, every nonzero prime ideal in $S_{c}$ has norm between $r$ and $q$.
In the following, we are going to verify that $K$ and $S_{c}$ satisfy the generalized Golod–Shafarevich condition (33) for $p=2$. According to Proposition 5, (38) will then be deduced immediately from (34).
For any given prime $p$, the $p$-rank $d_{p}Cl_{K}$ has a lower bound as follows (see Schoof [31] and Martinet [23]). Let $K/\mathbb{Q}$ be a Galois extension and $\theta$ denote the number of finite places of $\mathbb{Q}$ (i.e., primes of $\mathbb{Z}$) ramified in $K/\mathbb{Q}$. Assume that $\mathrm{Gal}(K/\mathbb{Q})$ is cyclic of order $p$. Then, we have $d_{p}Cl_{K}\geq\theta-1-\delta_{p}(\mathbb{Q})$, where $\delta_{p}(\mathbb{Q})$ equals $1$ if $\mathbb{Q}$ contains a primitive $p$th root of unity and 00 otherwise.
Taking $p=2$, we obtain $d_{2}Cl_{K}\geq\theta-2$ for the given quadratic field $K=\mathbb{Q}(\sqrt{\Delta_{K}})$. It is known that a finite place $p$ in $\mathbb{Q}$ is ramified in $K/\mathbb{Q}$ if and only if $p\mid\Delta_{K}$ (see [6, Proposition 3.4.3]). Thus, $\theta$ is equal to the number of primes which divide the discriminant $\Delta_{K}$ of $K$, i.e., $\theta=\ell$. Hence,
$d_{2}(Cl_{K})\geq\ell-2.$ (39)
The number of infinite places of $K=\mathbb{Q}(\sqrt{\Delta_{K}})$ is given by $|S_{\infty}|=2$ if $\Delta_{K}>0$ and by $|S_{\infty}|=1$ if $\Delta_{K}<0$. Thus,
$|S_{c}|+|S_{\infty}|\leq k+2.$ (40)
By (39), (40), and the given condition 2), we see that $K$ and $S_{c}$ satisfy the generalized Golod–Shafarevich condition (33) for $p=2$ as follows
$d_{2}(Cl_{K})\geq 2+2\sqrt{|S_{c}|+|S_{\infty}|+1}.$
Hence, (38) follows from (34) in Proposition 5. The proof is complete.  $\Box$
It is noticed that (38) in Theorem 38 is inferred from (34) in Proposition 5, which in turn is from (18) in Proposition 2.
In the following, we consider how to employ Theorem 38 for $q\in\mathbb{N}$ sufficiently large.
The lower bound for the rate function $R_{q}(\delta)$ given in Proposition 2, namely, the right-hand side of (18), is related to a positive integer $r$ with $2\leq r\leq q$. In order for this lower bound to be at least comparable to the Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ as given in (7), one cannot choose $r$ as small as $O(q^{1-\varepsilon})$ for some constant $\varepsilon\in(0,1)$.
Thus, we may choose $r=\lceil\lambda q\rceil$ with an arbitrary but fixed number $\lambda\in(0,1]$. If $\lambda=1$ is chosen, i.e., $r=q$, then, from Lemma 3, $A(q,q)\leq 1/\hat{c}$ for a constant $\hat{c}>0$. In this case, the right-hand side of (18) is not greater than $1-\delta-\frac{\hat{c}}{\log q}$. If we choose $r=\lceil\lambda q\rceil$ with $10<\lambda<$ 1, the right-hand side of (18) cannot exceed $1-\delta-\frac{\tilde{c}}{\log q}$ with $\tilde{c}=-(1-\delta)\log\lambda>0$ for $10<\delta<$ 1. Therefore, with a choice of $r=\lceil\lambda q\rceil$ for a constant $\lambda\in(0,1]$, the best attainable lower bound from (18) in Proposition 2 has asymptotically a logarithmic order in $q$ as the Gilbert–Varshamov bound in (7).
As indicated by (13) defining the number $\hat{N}_{r,q}(K)$, in [19] Lenstra employed an extended code from geometry of numbers, which takes infinite places into account, but it was chosen in [19] that $r=\lceil\frac{1}{2}q\rceil$ (i.e., $\lambda=\frac{1}{2}$) for nonzero prime ideals (i.e., finite places) of norm or its power between $r$ and $q$. By using the previous analysis in the same spirit, we can conclude that an achievable lower bound obtained by Lenstra in [19] for the rate function $R_{q}(\delta)$ at best has asymptotically a logarithmic order in $q$ as does the Gilbert–Varshamov bound in (7), even though the generalized Riemann hypothesis (GRH) was assumed in [19].
To beat the $q$-ary Gilbert–Varshamov bound $R_{\mathrm{GV}}(\delta,q)$ by a better lower bound for large $q$, we find that it is actually feasible with the help of Theorem 38. We can make use of (38) in Theorem 38 by choosing a sequence of numbers $\lambda_{q}$, satisfying $10<\lambda_{q}<$ 1 with $\lambda_{q}\to 1$ as $q\to\infty$, to define $r=r(q)=\lceil\lambda_{q}q\rceil$, while dealing with carefully the trade-off between choice of $r(q)$ (i.e., $\lambda_{q}$) and that of $\ell=\ell(q)$ and $k=k(q)$ to maximize the attainable lower bound from (38). This approach will need a subtle analysis of $k=k(q)$, given $r=r(q)$ and $\ell=\ell(q)$, in condition 3) of Theorem 38, and it relies on analytic number theory for the distribution of prime numbers.
In contrast to previous code constructions which have been based on a single infinite tower of number fields in [19] or of algebraic function fields over finite fields in [44], [35], [32], [9], and [3], the code construction in the proof of Theorem 38 uses two differentiated but complementary towers of Hilbert class fields with at least one being infinite.
It is ready now to establish proofs of Theorem 1 and Theorem 8, as will be presented in the subsequent sections. It is emphasized that the GRH is not assumed in the proofs, and each new lower bound in Theorem 1 and Theorem 8 is an unconditional result.
5 Proof of Theorem 1
In the following, we apply Theorem 38 for all $q\in\mathbb{N}$ with $q\geq Q\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\exp(29)$.
A suitable choice of the three positive integers $r,\ell,k\in\mathbb{N}$, depending on $q$, is given as follows. First, we choose
$\varepsilon_{q}=\frac{1}{(\log q)^{1/3}}.$ (41)
It is clear that $10<\varepsilon_{q}<$ 1 for all $q\geq Q$ and $\varepsilon_{q}\to 0$ as $q\to\infty$. We define
$r=r(q)=\lceil(1-\varepsilon_{q})^{2}q\rceil$ (42)
for $q\geq Q$. Next, we put
$x=\sqrt{q},\;\;\;y=(1-\varepsilon_{q})\sqrt{q}$
and choose
$\ell=\ell(q)=\lfloor x^{1/3}\rfloor$ (43)
for $q\geq Q$. Finally, we define
$k=k(q)=\left\lfloor\tfrac{1}{4}(\ell-2)^{2}-(\ell-2)\right\rfloor-2$ (44)
for $q\geq Q$.
We note that the three positive integers $r,\ell,k\in\mathbb{N}$, depending on $q$, as chosen in the above satisfy condition 1) and condition 2) of Theorem 38 for $q\geq Q$. We are going to prove that condition 3) of Theorem 38 also holds for the chosen $r,\ell,k\in\mathbb{N}$ for $q\geq Q$.
[…]
Proposition 6
If $x\geq 10^{3}$, then
$\left|\pi(x;4,3)-\frac{1}{2}\mathrm{Li}(x)\right|<0.53\frac{x}{(\log x)^{2}}.$ (45)
We are going to prove that the following cardinality of a subset
$N_{q}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left|\{p\in\mathbb{P}\mid r\leq p^{2}\leq q,\;p>p_{\ell},\mbox{ and }p\equiv 3\!\!\!\!\pmod{4}\}\right|$
satisfies $N_{q}\geq 2k$ for all $q\geq Q$.
It is computed that, for $q\geq Q=\exp(29)$, $\ell=\lfloor x^{1/3}\rfloor=\lfloor q^{1/6}\rfloor\geq\lfloor\exp(29/6)\rfloor=125$. Then, we can check that, for all $q\geq Q$,
$\sqrt{r}\geq(1-\varepsilon_{q})\sqrt{q}=(1-\varepsilon_{q})x>0.6745x>24x^{1/3}\log(x^{1/3})\geq 24\ell\log\ell>p_{\ell},$ (46)
where the last inequality follows from [1, Theorem 4.7].
Therefore, we get
$N_{q}=\left|\{p\in\mathbb{P}\mid r\leq p^{2}\leq q,\;p\equiv 3\!\!\!\!\pmod{4}\}\right|=\left|\{p\in\mathbb{P}\mid y^{2}\leq p^{2}\leq x^{2},\;p\equiv 3\!\!\!\!\pmod{4}\}\right|.$
Hence,
$N_{q}\geq\pi(x;4,3)-\pi(y;4,3)$
for all $q\geq Q$. According to Proposition 45, we obtain
$N_{q}\geq\pi(x;4,3)-\pi(y;4,3)\geq\frac{\mathrm{Li}(x)}{2}-\frac{\mathrm{Li}(y)}{2}-\frac{0.53x}{(\log x)^{2}}-\frac{0.53y}{(\log y)^{2}}\geq\frac{x-y}{2\log x}-\frac{1.06x}{(\log y)^{2}}$
where $x=\sqrt{q}$ and $y=(1-\varepsilon_{q})x$.
Now, we can verify that, for all $q\geq Q$,
$N_{q}\geq\frac{\varepsilon_{q}x}{2\log x}-\frac{1.06x}{(\log y)^{2}}\geq\frac{x}{2^{4/3}(\log x)^{4/3}}-\frac{1.06x}{(\log x+\log 0.6745)^{2}}\geq\frac{1}{2}x^{2/3}\geq\frac{1}{2}\ell^{2}\geq 2k.$
Therefore, the three positive integers $r,\ell,k\in\mathbb{N}$ as chosen in (42), (43), and (44), depending on $q$, satisfy all three conditions in Theorem 38 for all $q\geq Q$. Hence, it follows from (38) that
$R_{q}(\delta)\geq(1-\delta)\frac{\log r}{\log q}-\frac{\log\sqrt{D}}{k\log q}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}R_{\mathrm{NFC}}(\delta,q)$ (47)
for all $q\geq Q$ and $10<\delta<$ 1, where $D=4p_{1}\cdots p_{\ell}$.
Recall that the $q$-ary Gilbert–Varshamov bound is given by
$R_{\mathrm{GV}}(\delta,q)=1-\delta\log_{q}(q-1)-\delta\log_{q}\frac{1}{\delta}-(1-\delta)\log_{q}\frac{1}{1-\delta}$
for all $q\geq 2$ and $0<\delta<1-q^{-1}$. By (47), it suffices to verify that, for all $q\geq Q$,
$R_{\mathrm{NFC}}(1/2,q)>R_{\mathrm{GV}}(1/2,q),$ (48)
which implies that the $q$-ary Gilbert–Varshamov bound can be improved for all $q\geq Q$.
For all $q\geq Q$, we have
$R_{\mathrm{GV}}(1/2,q)=\frac{1}{2}+\left(\frac{1}{2}\log\frac{1}{1-q^{-1}}-\log 2\right)\frac{1}{\log q}<\frac{1}{2}-0.6839\frac{1}{\log q}.$ (49)
On the other hand, noting that
$\frac{1}{2}\log\frac{q}{r}\leq\frac{1}{2}\log\frac{q}{(1-\varepsilon_{q})^{2}q}=\log\frac{1}{1-\varepsilon_{q}}<\log\frac{1}{0.6745}<0.3938$
for all $q\geq Q$, it follows from (47) that
$R_{\mathrm{NFC}}(1/2,q)=\frac{1}{2}-\left(\frac{1}{2}\log\frac{q}{r}+\frac{\log D}{2k}\right)\frac{1}{\log q}>\frac{1}{2}-\left(0.3938+\frac{\log D}{2k}\right)\frac{1}{\log q}$ (50)
for all $q\geq Q$.
Therefore, in order to prove (48) for all $q\geq Q$, it suffices to verify that, for all $q\geq Q$,
$\frac{\log D}{2k}\leq 0.6839-0.3938=0.2901$ (51)
where $D=4p_{1}\cdots p_{\ell}$ and $k=\left\lfloor\tfrac{1}{4}(\ell-2)^{2}-(\ell-2)\right\rfloor-2$.
In terms of Chebyshev’s $\vartheta$-function [1, p. 75] defined by $\vartheta(x)=\sum_{p\leq x}\log p$ for $x\geq 2$, the verification of (51) for all $q\geq Q$ is equivalent to proving
$\vartheta(p_{\ell})\leq-\log 4+0.5802\left[\left\lfloor\tfrac{1}{4}(\ell-2)^{2}-(\ell-2)\right\rfloor-2\right]$ (52)
for all $q\geq Q$.
Since $q\geq Q=\exp(29)$, we have $\ell=\lfloor q^{1/6}\rfloor\geq 125$ and hence $p_{\ell}\geq p_{125}=691$. Thus, by applying an explicit bound for $\vartheta$-function given in [30, p. 211], we have
$\vartheta(p_{\ell})<\left(1+\frac{3}{\log p_{\ell}}\right)p_{\ell}.$ (53)
Moreover, by using an explicit bound given in [4, p. 433], we have
$\frac{p_{\ell}}{\log p_{\ell}}<\pi(p_{\ell})=\ell.$ (54)
Noting the last inequality in (46), i.e., $p_{\ell}<24\ell\log\ell$, it follows from (54) that
$p_{\ell}<\ell\log p_{\ell}<\ell\log\ell+\ell\log\log p_{\ell}<\ell\log\ell+\ell\log\log\ell^{2}.$
Hence, by using (53) and (54), we get
$\vartheta(p_{\ell})<3\ell+p_{\ell}<(3+\log 2)\ell+\ell\log\ell+\ell\log\log\ell.$
Therefore, in order to prove (52) for all $q\geq Q$, it is sufficient to verify that
$3.7\ell+\ell\log\ell+\ell\log\log\ell\leq-1.39+0.58\left[\tfrac{1}{4}(\ell-2)^{2}-(\ell-2)-3\right]$
for all $\ell\geq 125$. This can be easily checked. The proof of Theorem 1 is complete.  $\Box$
6 Proof of Theorem 8
The upper bound in (8), namely $\eta(\delta)\leq 1$ for any real number $\delta\in(0,1)$, is deduced from the definition of the growth $\eta(\delta)$ in (6) and the Plotkin bound for rate function $R_{q}(\delta)$ in (5). In the following, we are going to establish the nontrivial lower bound in (8).
We apply Theorem 38 for $q\in\mathbb{N}$ sufficiently large, which needs an appropriate choice of the three positive integers $r,\ell,k\in\mathbb{N}$, depending on $q$.
First, we choose
$\varepsilon_{q}=\frac{(\log q)\log\log q}{C_{0}q^{1/6}}$ (55)
where $C_{0}>1$ is a suitable absolute constant such that $10<\varepsilon_{q}<$ 1 for all $q\geq 27$. It is seen that $\varepsilon_{q}\to 0$ as $q\to\infty$. We define
$r=r(q)=\lceil(1-\varepsilon_{q})^{2}q\rceil$ (56)
for $q\geq 27$.
Next, we set
$x=\sqrt{q},\;\;\;y=(1-\varepsilon_{q})\sqrt{q}$
and choose
$\ell=\ell(q)=\lfloor x^{1/3}\rfloor$ (57)
for $q\geq 27$.
Finally, we define
$k=k(q)=\left\lfloor\tfrac{1}{4}(\ell-2)^{2}-(\ell-2)\right\rfloor-2$ (58)
for $q\in\mathbb{N}$ such that $\ell(q)\geq 10$.
The three positive integers $r,\ell,k\in\mathbb{N}$, depending on $q$, as chosen in the above satisfy condition 1) and condition 2) of Theorem 38 for $q$ large enough. In the following, we need to check that condition 3) of Theorem 38 is also satisfied by the chosen $r,\ell,k\in\mathbb{N}$ provided that $q$ is sufficiently large.
We are going to prove that the following cardinality of a subset
$N_{q}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left|\{p\in\mathbb{P}\mid\sqrt{r}\leq p\leq\sqrt{q},\;p>p_{\ell},\mbox{ and }p\equiv 3\!\!\!\!\pmod{4}\}\right|$
satisfies $N_{q}\geq 2k$ for $q$ large enough.
Noting that $\ell=\lfloor x^{1/3}\rfloor=\lfloor q^{1/6}\rfloor$, we get
$\sqrt{r}\geq(1-\varepsilon_{q})\sqrt{q}=(1-\varepsilon_{q})x>0.99x>2x^{1/3}\log(x^{1/3})\geq 2\ell\log\ell>p_{\ell}$
for $q$ large enough.
Therefore, we have
$N_{q}=\left|\{p\in\mathbb{P}\mid\sqrt{r}\leq p\leq\sqrt{q},\;p\equiv 3\!\!\!\!\pmod{4}\}\right|=\left|\{p\in\mathbb{P}\mid y\leq p\leq x,\;p\equiv 3\!\!\!\!\pmod{4}\}\right|.$
Hence, for $q$ sufficiently large,
$N_{q}\geq\pi(x;4,3)-\pi(y;4,3)$ (59)
where $x=\sqrt{q}$ and $y=(1-\varepsilon_{q})x$.
For the given functions $f(q)$ and $g(q)$ of $q\in\mathbb{N}$, if $\lim_{q\to\infty}\frac{f(q)}{g(q)}=1$, we say that the two functions $f(q)$ and $g(q)$ are asymptotically equivalent, denoted by $f(q)\sim g(q)$, as $q\to\infty$. By the prime number theorem, we know that the positive integer $D=4p_{1}\cdots p_{\ell}$ defined in Theorem 38 satisfies
$\log D=\log 4+\sum_{p\leq p_{\ell}}\log p\sim p_{\ell}\sim\ell\log\ell$ (60)
as $q\to\infty$ (see [1, Chapter 4]). By (57), we have $\ell\sim q^{1/6}$ and hence $\log\ell\sim\frac{1}{6}\log q$ and $p_{\ell}\sim\ell\log\ell\sim\frac{1}{6}q^{1/6}\log q$ as $q\to\infty$.
We define $h=x-y$ with $x=\sqrt{q}$ and $y=(1-\varepsilon_{q})x$, which satisfies
$h=\varepsilon_{q}x\sim C_{1}x^{2/3}(\log x)\log\log x$ (61)
for $q$ sufficiently large with a suitable constant $C_{1}>0$.
By using theory on the distribution of primes in arithmetic progressions and in short intervals (see, e.g., [2, Theorem 3], [25, p. 656, Theorem]), we know that, for any fixed real number $C_{2}\in(0,1)$,
$\pi(x;4,3)-\pi(y;4,3)=\pi(x;4,3)-\pi(x-h;4,3)\geq C_{2}\frac{h}{2\log x}$ (62)
for $x=\sqrt{q}$ sufficiently large. Note that (62) holds for all $h$ with $x^{\beta}\leq h\leq x(\log x)^{-1}$ where $\beta\in\mathbb{R}$ is any constant satisfying $15/8<\beta<$ 1, and it is a generalization of Ingham’s theorem on the distribution of primes in short intervals [18].
It follows from (59), (62), and (61) that
$N_{q}\geq C_{2}\frac{h}{2\log x}\sim C_{3}x^{2/3}\log\log x$
for $q$ sufficiently large with a suitable constant $C_{3}>0$. On the other hand, we have
$2k\sim\frac{1}{2}\ell^{2}\sim\frac{1}{2}x^{2/3}$
for $q$ sufficiently large. Hence, $N_{q}\geq 2k$ for $q$ large enough.
In summary, we have shown that the three positive integers $r,\ell,k\in\mathbb{N}$ as chosen in (56), (57), and (58), depending on $q$, satisfy all three conditions in Theorem 38 for $q\in\mathbb{N}$ sufficiently large.
Now, by using (38), we obtain
$R_{q}(\delta)\geq(1-\delta)\frac{\log r}{\log q}-\frac{\log\sqrt{D}}{k\log q}=1-\delta-(1-\delta)\frac{\log(q/r)}{\log q}-\frac{\log D}{2k\log q}$ (63)
for $q\geq C_{4}$ and $10<\delta<$ 1, where $C_{4}>0$ is an absolute constant.
It is clear from (56) that
$\log\frac{q}{r}\leq\log\frac{q}{(1-\varepsilon_{q})^{2}q}=2\log\frac{1}{1-\varepsilon_{q}}<3\varepsilon_{q}$ (64)
for $q$ sufficiently large.
Moreover, we can see that
$\frac{\log D}{4k}\sim\frac{\log D}{\ell^{2}}\sim\frac{\log\ell}{\ell}\sim\frac{\log x}{3x^{1/3}}\sim\frac{C_{0}\varepsilon_{q}}{C_{5}\log\log q}$
as $q\to\infty$, where $C_{5}>0$ is some constant.
Hence, for $q$ large enough,
$\frac{\log D}{2k}<2C_{0}\varepsilon_{q}.$ (65)
It follows from (63), (64), and (65) that there exists an absolute constant $C_{6}>0$ such that, for all $q\geq C_{6}$ and $10<\delta<$ 1,
$R_{q}(\delta)>1-\delta-\frac{3\varepsilon_{q}}{\log q}-\frac{2C_{0}\varepsilon_{q}}{\log q}>1-\delta-\frac{5C_{0}\varepsilon_{q}}{\log q}=1-\delta-5\frac{\log\log q}{q^{1/6}}.$
Consequently, there exists an absolute constant $C>5$ such that
$R_{q}(\delta)>1-\delta-C\frac{\log\log q}{q^{1/6}}$
for all $q\geq 27$ and $10<\delta<$ 1. Hence,
$\eta(\delta)=\liminf_{q\rightarrow\infty}\frac{1}{\log q}\log[1-\delta-R_{q}(\delta)]^{-1}\geq\frac{1}{6}$
for $10<\delta<$ 1.
Therefore, the proof of Theorem 8 is complete.  $\Box$